# A Simulation Based Inference Approach to the Dynamics of the MW–LMC System - Validation

### v1

@@ 2.2 The Milky Way Stellar Halo
To generate a mock MW stellar halo for each simulation, we draw phase-space samples from radially biased distribution functions as implemented in agama (Vasiliev, 2019). This requires instances of a tracer density profile, a potential and a prescription for the radial velocity anisotropy. We use a Dehnen tracer density profile (Dehnen, 1993) and an NFW profile for the potential (Navarro et al., 1996, 1997) with relevant parameters adopted from each unique MW–LMC simulation; see Sec. 2.3. We implement the radial bias of stellar velocities (Binney & Tremaine, 2008) using the velocity anisotropy parameter as implemented in agama (Vasiliev, 2019):
@@
where $\beta_{0}$ is the limiting value of anisotropy in the centre, and if $r_{a}<\infty$, the anisotropy coefficient tends to 1 at large $r$ (Osipkov–Merritt profile), otherwise it is constant and equal to $\beta_{0}$ over all scales which we adopt. Han et al. (2024) found that $\beta$ varies as a function of distance and on-sky angle. However, other studies have found that $\beta$ takes an almost constant value throughout the Galaxy (e.g., Bird et al., 2021; Cunningham et al., 2019; Chandra et al., 2025). We sample the distribution function to return a set of $4500$ phase-space coordinates to represent the MW stellar halo at some early time, i.e. well before the LMC has caused any disequilibrium in the MW. We adopt this number of particles for the stellar halo as this approximately matches the average number density of stars beyond $50\,\rm kpc$ in both the Dark Energy Spectroscopic Instrument (DESI) Blue Horizontal Branch (BHB, Byström et al., 2024) and the combined H3, + SEGUE + MagE red giant samples (Chandra et al., 2025). For a given MW–LMC potential with reflex motion, we integrate all particles in the stellar halo to present-day over the last $2.2\,\mathrm{Gyr}$. We choose the amount of time to evolve the MW–LMC system as it matches the live simulation time in Garavito-Camargo et al. (2019).
From the final distribution of stellar halo particles, we measure the reflex motion of the MW in response to the infalling LMC. To do this, we adapt the method from Petersen & Peñarrubia (2021) which has subsequently been used in other works (e.g., Yaaqib et al., 2024; Chandra et al., 2025). We fit an on-sky velocity model that contains nine free parameters. We model the dipole reflex motion using Galactocentric Cartesian velocities ($v_{x},v_{y},v_{z}$). This is subtly distinct from using the magnitude of reflex motion, $v_{\mathrm{travel}}$, and its direction on the sphere, $(l_{\mathrm{apex}},b_{\mathrm{apex}})$, directly as parameters (Petersen & Peñarrubia, 2021). We do this for simplicity and to avoid inefficient convergence during the maximum likelihood estimates of these parameters when using the mock stellar halo data. The Galactocentric velocities are easily converted to their spherical equivalents by:
@@
Where $\vec{v}$ is the mean Galactocentric halo velocity vector. We project the velocities to observable coordinates using equs. 3 & 4 in Petersen & Peñarrubia (2021). To find the maximum likelihood estimates for these parameters given each mock stellar halo data, we minimise a Gaussian log-likelihood for the 1-dimensional line of sight velocities and 2-dimensional proper motions using scipy.optimize (see equs. 6 & 8 in Petersen & Peñarrubia, 2021). We return the maximum likelihood estimates for all of the reflex motion model parameters. However, in the context of this work, we will only comment on $v_{\mathrm{travel}},l_{\mathrm{apex}},\rm{and}\,b_{\mathrm{apex}}$.
@@ 2.3 Simulation Priors
We summarise the prior probability distributions used for free model parameters in Table. 2.3. The first three parameters are the MW mass, the infall LMC mass [^2] Note, as the LMC mass is fixed throughout each rigid simulation, the infall and present-day LMC masses are the same., and dynamical friction strength, which are our target inference parameters. The middle set of parameters describe the LMC present-day position and velocity with their distributions inspired by the values in sec. 3.1 and table 2 of Vasiliev (2023). The final set of parameters are the anisotropy parameter, $\beta_{0}$, and the Dehnen tracer density profile scale length, $r_{\rm Dehnen}$, that initialise the mock MW stellar haloes. In total, we run $128,000$ MW–LMC simulations with unique initial conditions for the MW and LMC as described above.
@@ 3 Simulation Based Inference
The simplest form of SBI is known as Approximate Bayesian Computation (ABC, e.g., Rubin, 1984; Pritchard et al., 1999; Fearnhead & Prangle, 2010). The ABC framework selects forward simulations that are the most similar to the observed data based on some distance measure involving the summary statistics of the simulation. Another way to compute the posterior is via Density Estimation Likelihood Free Inference (DELFI). In this approach, forward simulations are used to learn a conditional density distribution of the data $D_{\text{obs}}$, given the simulation parameters $\theta$, using a density estimation algorithm, e.g., normalising flows, that utilise a series of transformations to convert a simple base distribution into the desired probability distribution (Jimenez Rezende & Mohamed, 2015). We use the package sbi (Tejero-Cantero et al., 2020), and estimate the posterior distribution from the forward simulations using Masked Autoregressive Flows (MAF, Papamakarios et al., 2017, 2019) with $5$ layers in the neural network, each with $50$ nodes. DELFI is advantageous over the simpler ABC approach as it does not rely on a choice of a distance measure and it uses all available forward simulations to build the posterior distribution, making it far more efficient (Alsing et al., 2019). Additionally, once a normalising flow has been trained on a precomputed simulation dataset, the posterior can be returned for many observations without having to retrain the flow. This is known as amortisation (Mittal et al., 2025) and is a benefit over more classical approaches to estimate the posterior such as Markov Chain Monte Carlo as the chains need to be re-evaluated every time a new observation becomes available.
Often, some form of data compression is required (e.g., Alsing et al., 2018; Alsing & Wandelt, 2019; Heavens et al., 2020; Jeffrey et al., 2021; Widmark & Johnston, 2025). However, our problem has only a few parameters of interest and data points, so no data compression is required. We proceed by generating a large number of rigid MW–LMC simulations that cover broad regions of parameter space. Generally, for SBI, the more simulations that are available to use, the better. Recently, Bairagi et al. (2025) estimated the number of simulations that are required, $\sim\mathcal{O}(10^{3}-10^{4})$, when using SBI within a cosmological context. The simulations used in this work are described in Sec. 2. We will then use a MAF density estimator from the sbi package (Tejero-Cantero et al., 2020) to directly obtain the posterior distribution that can be evaluated at any observed data point for any data realisation, i.e, $p(\theta|D_{\text{obs}},I)$. In Sec. 5, we provide a complete discussion of posterior probability diagnostic checks.
@@ 4 Results
We perform a series of tests to assess the validity of our SBI framework. First, in Sec. 4.1 , we select a random simulation from our large set of rigid MW–LMC simulations. In this closed-loop test, we aim to recover some of its true simulation parameters by adopting some of the other simulation parameters as ‘data points’. Second, in Sec. 4.2, we extend this concept to recover simulation parameters from well-known idealised deforming MW–LMC models (Garavito-Camargo et al., 2019, 2021b). Finally, in Sec. 4.3, we push this formalism further to a MW–LMC analogue pair identified in the FIRE latte suite of (Wetzel et al., 2016) cosmological simulations.
@@ 4.1 On rigid MW–LMC models
To begin with a simple example, we aim to recover true parameters of a rigid MW–LMC simulation that uses our prescription for the potential as outlined in Sec. 2. By selecting a simulation from our sample, we expect to be able to find posterior distributions which enclose the true parameters of the simulation. The randomly chosen simulation has parameter values as follows: The LMC has a present-day Galactocentric position, $\mathbf{x}_{\mathrm{LMC}}=\{-1.1,-28.0,-21.5\}\,\rm kpc$, and velocity, $\mathbf{v}_{\mathrm{LMC}}=\{-87.1,-198.9,135.4\}\,\rm km\,\rm s^{-1}$. The LMC mass is $\rm M_{LMC}=23\times 10^{10}\,\rm M_{\odot}$, with scale radius, $a_{\rm LMC}=18.9\,\rm kpc$. The MW mass is $\rm M_{200,MW}=9.1\times 10^{11}\,\rm M_{\odot}$, with concentration $c_{200}=14.5$. For the stellar halo, the anisotropy parameter is $\beta_{0}=0.76$, and the Dehnen scale radius is $11.5\,\rm kpc$ for the density profile. In this MW–LMC potential, the MW reflex parameters as measured from the stellar halo are $v_{\mathrm{travel}}=64.5\,\rm km\,\rm s^{-1},l_{\mathrm{apex}}=39.1^{\circ},b_{\mathrm{apex}}=-67.8^{\circ}$; see Sec. 2.2. The dynamical friction scalar strength used is $\lambda_{\rm DF}=0.7$.
@@
In Fig. 1 we show the joint, and individual, posterior distributions for the LMC and MW total masses using a set of $128,000$ MW–LMC simulations. The true values for the MW mass, LMC mass, and dynamical friction strength (yellow square and lines) all lie within the $16^{\rm th}$ and $84^{\rm th}$ percentiles of the posterior (purple shaded region in 1D insets). We additionally show the prior distributions in grey to give intuition about the constraining power in each parameter. The LMC mass and dynamical friction strength are both accurately and precisely constrained by the supplied data values for the LMC present-day position, velocity, and the reflex motion parameters. On the other hand, the MW mass is accurately constrained, although its spread, i.e., its precision, remains similar to the original prior. This suggests that the learned MW mass posterior is somewhat insensitive to the given data points. Much of the constraining power for the LMC comes from providing the data points for $v_{\rm travel},l_{\rm apex},b_{\rm apex}$; see the all-sky projection for the direction of the apex on-sky (pink cross). This is not unexpected as the LMC mass and the magnitude of the travel velocity, $v_{\rm travel}$, are positively correlated (e.g., Petersen & Peñarrubia, 2021). This implies that providing $v_{\rm travel}$ as a data point is crucial in breaking degeneracies for the returned posteriors.
@@ 4.2 On deforming MW–LMC models
Figure 2: Validation on a deforming MW–LMC models: The joint, and individual, posterior distributions for MW mass, LMC mass and dynamical friction strength when there is knowledge of the LMC present day position, velocity and the reflex motion parameters (purple). We also show the parameter prior distributions in grey. For the 1D posterior panels we show the $16^{\rm th}$ and $84^{\rm th}$ percentiles as shaded regions. The yellow squares/lines represent the true values from the fiducial deforming simulation of Garavito-Camargo et al. (2019). The reflex motion parameters $v_{\mathrm{travel}},l_{\mathrm{apex}},b_{\mathrm{apex}}$ break model parameter degeneracies to add constraining power, particularly in the LMC mass. The all-sky projection displays the average solar corrected radial velocity, $v_{\rm GSR}$, for all stellar halo particles beyond $50\,\rm kpc$. The net signal across the entire sky is subtracted. The pink cross denotes the reflex motion apex for this simulation. The past orbit of the LMC in this potential is shown as the black dashed line. Orbits computed using posterior samples for the masses and dynamical friction strength, for fixed LMC present-day conditions, are shown as the grey shaded region.
We assess the validity of our sample of rigid MW–LMC simulations to be able to confidently capture the dynamics of the more complicated deforming MW–LMC simulations. In other words, we want to assess whether the rigid set-up can be upscaled in their fidelity to resemble the more complex, deforming MW–LMC simulations. In particular, we aim to recover the true parameters from the Garavito-Camargo et al. (2019) deforming MW–LMC simulations. We choose to recover parameters from the fiducial simulation run. In this simulation, the LMC has a present-day Galactocentric position, $\mathbf{x}_{\mathrm{LMC}}=\{-2.32,-45.57,-21.45\}\,\rm kpc$ and velocity $\mathbf{v}_{\mathrm{LMC}}=\{-59.9,-249.3,227.9\}\,\rm km\,\rm s^{-1}$. The fiducial infall LMC mass is $M_{\rm LMC}=18\times 10^{10}\,\rm M_{\odot}$ with scale radius $a_{\rm LMC}=20\,\rm kpc$. The fiducial MW mass is $M_{\rm 200,MW}=10.3\times 10^{11}\,\rm M_{\odot}$ with concentration $c_{200}=11.0$. To measure the MW reflex parameters, we take all of the MW halo dark matter particles from the simulation and apply the same particle selection criteria as in Sec. 2.2 for our mock stellar haloes. This ensures that the number of particles approximately matches the average number density of stars beyond $50\,\rm kpc$ in the DESI BHB sample within the DESI sky-coverage footprint (Byström et al., 2024, adopting equ. 11 for the DESI footprint). The resulting reflex motion parameters as measured from this simulation are $v_{\mathrm{travel}}=27.9\,\mathrm{km}\,\mathrm{s}^{-1},l_{\mathrm{apex}}=-49.1^{\circ},b_{\mathrm{apex}}=-80.4^{\circ}$; see the all-sky projection in Fig. 2 for the direction of the apex on-sky (pink cross).
@@
[…]
Figure 3: Comparison of the true and inferred infall LMC masses: We compare the simulation truth and returned SBI posterior LMC mass for all of the MW–LMC deforming simulations in Garavito-Camargo et al. (2019). We show the $16^{\rm th}-84^{\rm th}$ percentiles of the LMC mass posteriors as error bars. For all deforming MW–LMC simulations except the most massive LMC, the true and SBI returned LMC masses are consistent, i.e., they overlap on the 1:1 dashed grey line, within the $1\,\sigma$ confidence interval. Additionally, we show the re-weighted SBI masses, which account for any biases in the choice of the LMC mass prior, as the fainter points and errors. We colour each point by the measured simulation travel velocity to demonstrate its expected increase with increasing LMC mass.

### v2

@@ 2.2 The Milky Way Stellar Halo
To generate a mock MW stellar halo for each simulation, we draw phase-space samples from radially biased distribution functions as implemented in agama (2019MNRAS.482.1525V). This requires instances of a tracer density profile, a potential and a prescription for the radial velocity anisotropy. We use a Dehnen tracer density profile (Dehnen1993) and an NFW profile for the potential (Navarro1996; Navarro1997) with relevant parameters adopted from each unique MW–LMC simulation; see Sec. 2.3. We implement the radial bias of stellar velocities (Binney2008) using the velocity anisotropy parameter as implemented in agama (2019MNRAS.482.1525V):
@@
where $\beta_{0}$ is the limiting value of anisotropy in the centre, and if $r_{a}<\infty$, the anisotropy coefficient tends to 1 at large $r$ (Osipkov–Merritt profile), otherwise it is constant and equal to $\beta_{0}$ over all scales which we adopt. Han2024 found that $\beta$ varies as a function of distance and on-sky angle. However, other studies have found that $\beta$ takes an almost constant value throughout the Galaxy (e.g., Bird2021; Cunningham2019; Chandra2025). We sample the distribution function to return a set of $4500$ phase-space coordinates to represent the MW stellar halo at some early time, i.e. well before the LMC has caused any disequilibrium in the MW. We adopt this number of particles for the stellar halo as this approximately matches the average number density of stars beyond $50\,\rm kpc$ in both the Dark Energy Spectroscopic Instrument (DESI) Blue Horizontal Branch (BHB, Bystrom2025) and the combined H3, + SEGUE + MagE red giant samples (Chandra2025). For a given MW–LMC potential with reflex motion, we integrate all particles in the stellar halo to present-day over the last $2.2\,\mathrm{Gyr}$. We choose the amount of time to evolve the MW–LMC system as it matches the live simulation time in Garavito-Camargo2019.
From the final distribution of stellar halo particles, we measure the reflex motion of the MW in response to the infalling LMC. To do this, we adapt the method from Petersen2021 which has subsequently been used in other works (e.g., Yaaqib2024; Chandra2025). We fit an on-sky velocity model that contains nine free parameters. We model the dipole reflex motion using Galactocentric Cartesian velocities ($v_{x},v_{y},v_{z}$). This is subtly distinct from using the magnitude of reflex motion, $v_{\mathrm{travel}}$, and its direction on the sphere, $(l_{\mathrm{apex}},b_{\mathrm{apex}})$, directly as parameters (Petersen2021). We do this for simplicity and to avoid inefficient convergence during the maximum likelihood estimates of these parameters when using the mock stellar halo data. The Galactocentric velocities are easily converted to their spherical equivalents by:
@@
Where $\vec{v}$ is the mean Galactocentric halo velocity vector. We project the velocities to observable coordinates using equs. 3 & 4 in Petersen2021. To find the maximum likelihood estimates for these parameters given each mock stellar halo data, we minimise a Gaussian log-likelihood for the 1-dimensional line of sight velocities and 2-dimensional proper motions using scipy.optimize (see equs. 6 & 8 in Petersen2021). We return the maximum likelihood estimates for all of the reflex motion model parameters. However, in the context of this work, we will only comment on $v_{\mathrm{travel}},l_{\mathrm{apex}},\rm{and}\,b_{\mathrm{apex}}$.
@@ 2.3 Simulation Priors
We summarise the prior probability distributions used for free model parameters in Table. 2.3. The first three parameters are the MW mass, the infall LMC mass [^2] Note, as the LMC mass is fixed throughout each rigid simulation, the infall and present-day LMC masses are the same., and dynamical friction strength, which are our target inference parameters. The middle set of parameters describe the LMC present-day position and velocity with their distributions inspired by the values in sec. 3.1 and table 2 of Vasiliev2023. The final set of parameters are the anisotropy parameter, $\beta_{0}$, and the Dehnen tracer density profile scale length, $r_{\rm Dehnen}$, that initialise the mock MW stellar haloes. In total, we run $128,000$ MW–LMC simulations with unique initial conditions for the MW and LMC as described above.
@@ 3 Simulation Based Inference
The simplest form of SBI is known as Approximate Bayesian Computation (ABC, e.g., Rubin1984; Pritchard1999; Fearnhead2010). The ABC framework selects forward simulations that are the most similar to the observed data based on some distance measure involving the summary statistics of the simulation. Another way to compute the posterior is via Density Estimation Likelihood Free Inference (DELFI). In this approach, forward simulations are used to learn a conditional density distribution of the data $D_{\text{obs}}$, given the simulation parameters $\theta$, using a density estimation algorithm, e.g., normalising flows, that utilise a series of transformations to convert a simple base distribution into the desired probability distribution (JimenezRezende2015). We use the package sbi (tejero-cantero2020sbi), and estimate the posterior distribution from the forward simulations using Masked Autoregressive Flows (MAF, Papamakarios2017; Papamakarios2019) with $5$ layers in the neural network, each with $50$ nodes. DELFI is advantageous over the simpler ABC approach as it does not rely on a choice of a distance measure and it uses all available forward simulations to build the posterior distribution, making it far more efficient (Alsing2019). Additionally, once a normalising flow has been trained on a precomputed simulation dataset, the posterior can be returned for many observations without having to retrain the flow. This is known as amortisation (Mittal2025) and is a benefit over more classical approaches to estimate the posterior such as Markov Chain Monte Carlo as the chains need to be re-evaluated every time a new observation becomes available.
Often, some form of data compression is required (e.g., Alsing2018; Alsing2019b; Heavens2020; Jeffrey2021; Widmark2025). However, our problem has only a few parameters of interest and data points, so no data compression is required. We proceed by generating a large number of rigid MW–LMC simulations that cover broad regions of parameter space. Generally, for SBI, the more simulations that are available to use, the better. Recently, Bairagi2025 estimated the number of simulations that are required, $\sim\mathcal{O}(10^{3}-10^{4})$, when using SBI within a cosmological context. The simulations used in this work are described in Sec. 2. We will then use a MAF density estimator from the sbi package (tejero-cantero2020sbi) to directly obtain the posterior distribution that can be evaluated at any observed data point for any data realisation, i.e, $p(\theta|D_{\text{obs}},I)$. In Sec. 5, we provide a complete discussion of posterior probability diagnostic checks.
@@ 4 Results
We perform a series of tests to assess the validity of our SBI framework. First, in Sec. 4.1 , we select a random simulation from our large set of rigid MW–LMC simulations. In this closed-loop test, we aim to recover some of its true simulation parameters by adopting some of the other simulation parameters as ‘data points’. Second, in Sec. 4.2, we extend this concept to recover simulation parameters from well-known idealised deforming MW–LMC models (Garavito-Camargo2019; Garavito-Camargo2021). Finally, in Sec. 4.3, we push this formalism further to a MW–LMC analogue pair identified in the FIRE latte suite of (Wetzel2016) cosmological simulations. In all figures we use $10,000$ samples drawn from their respective posterior density distributions.
@@ 4.1 On rigid MW–LMC models
To begin with a simple example, we aim to recover true parameters of a rigid MW–LMC simulation that uses our prescription for the potential as outlined in Sec. 2. By selecting a simulation from our sample, which is not used to train the SBI framework, we expect to be able to find posterior distributions which enclose the true parameters of the simulation. The randomly chosen simulation has parameter values as follows: The LMC has a present-day Galactocentric position, $\mathbf{x}_{\mathrm{LMC}}=\{-1.1,-28.0,-21.5\}\,\rm kpc$, and velocity, $\mathbf{v}_{\mathrm{LMC}}=\{-87.1,-198.9,135.4\}\,\rm km\,\rm s^{-1}$. The LMC mass is $\rm M_{LMC}=23\times 10^{10}\,\rm M_{\odot}$, with scale radius, $a_{\rm LMC}=18.9\,\rm kpc$. The MW mass is $\rm M_{200,MW}=9.1\times 10^{11}\,\rm M_{\odot}$, with concentration $c_{200}=14.5$. For the stellar halo, the anisotropy parameter is $\beta_{0}=0.76$, and the Dehnen scale radius is $11.5\,\rm kpc$ for the density profile. In this MW–LMC potential, the MW reflex parameters as measured from the stellar halo are $v_{\mathrm{travel}}=64.5\,\rm km\,\rm s^{-1},l_{\mathrm{apex}}=39.1^{\circ},b_{\mathrm{apex}}=-67.8^{\circ}$; see Sec. 2.2. The dynamical friction scalar strength used is $\lambda_{\rm DF}=0.7$.
@@
In Fig. 1 we show the joint, and individual, posterior distributions for the LMC and MW total masses using a set of $128,000$ MW–LMC simulations. The true values for the MW mass, LMC mass, and dynamical friction strength (yellow square and lines) all lie within the $16^{\rm th}$ and $84^{\rm th}$ percentiles of the posterior (purple shaded region in 1D insets). We additionally show the non-normalised prior distributions in grey to give intuition about the constraining power in each parameter. The LMC mass and dynamical friction strength are both accurately and precisely constrained by the supplied data values for the LMC present-day position, velocity, and the reflex motion parameters. On the other hand, the MW mass is accurately constrained, although its spread, i.e., its precision, remains similar to the original prior. This suggests that the learned MW mass posterior is somewhat insensitive to the given data points. Much of the constraining power for the LMC comes from providing the data points for $v_{\rm travel},l_{\rm apex},b_{\rm apex}$; see the all-sky projection for the direction of the apex on-sky (pink cross). This is not unexpected as the LMC mass and the magnitude of the travel velocity, $v_{\rm travel}$, are positively correlated (e.g., Petersen2021). This implies that providing $v_{\rm travel}$ as a data point is crucial in breaking degeneracies for the returned posteriors. Indeed, we have performed inference on the same model parameters but only conditioning on the LMC present day position and velocity. In this case, we find that the posterior closely resembles the prior, implying that most of the constraining power comes from the reflex motion measurements.
@@ 4.2 On deforming MW–LMC models
Figure 2: Validation on a deforming MW–LMC models: The joint, and individual, posterior distributions for MW mass, LMC mass and dynamical friction strength when there is knowledge of the LMC present day position, velocity and the reflex motion parameters (purple). We also show the parameter prior distributions in grey. For the 1D posterior panels we show the $16^{\rm th}$ and $84^{\rm th}$ percentiles as shaded regions. The yellow squares/lines represent the true values from the fiducial deforming simulation of Garavito-Camargo2019. The reflex motion parameters $v_{\mathrm{travel}},l_{\mathrm{apex}},b_{\mathrm{apex}}$ break model parameter degeneracies to add constraining power, particularly in the LMC mass. The all-sky projection displays the average solar corrected radial velocity, $v_{\rm GSR}$, for all stellar halo particles beyond $50\,\rm kpc$. The net signal across the entire sky is subtracted. The pink cross denotes the reflex motion apex for this simulation. The past orbit of the LMC in this potential is shown as the black dashed line. Orbits computed using posterior samples for the masses and dynamical friction strength, for fixed LMC present-day conditions, are shown as the grey shaded region.
We assess the validity of our sample of rigid MW–LMC simulations to be able to confidently capture the dynamics of the more complicated deforming MW–LMC simulations. In other words, we want to assess whether the rigid set-up can be upscaled in their fidelity to resemble the more complex, deforming MW–LMC simulations. In particular, we aim to recover the true parameters from the Garavito-Camargo2019 deforming MW–LMC simulations. We choose to recover parameters from the fiducial simulation run. In this simulation, the LMC has a present-day Galactocentric position, $\mathbf{x}_{\mathrm{LMC}}=\{-2.32,-45.57,-21.45\}\,\rm kpc$ and velocity $\mathbf{v}_{\mathrm{LMC}}=\{-59.9,-249.3,227.9\}\,\rm km\,\rm s^{-1}$. The fiducial infall LMC mass is $M_{\rm LMC}=18\times 10^{10}\,\rm M_{\odot}$ with scale radius $a_{\rm LMC}=20\,\rm kpc$. The fiducial MW mass is $M_{\rm 200,MW}=10.3\times 10^{11}\,\rm M_{\odot}$ with concentration $c_{200}=11.0$. To measure the MW reflex parameters, we take all of the MW halo dark matter particles from the simulation and apply the same particle selection criteria as in Sec. 2.2 for our mock stellar haloes. This ensures that the number of particles approximately matches the average number density of stars beyond $50\,\rm kpc$ in the DESI BHB sample within the DESI sky-coverage footprint (Bystrom2025, adopting equ. 11 for the DESI footprint). The resulting reflex motion parameters as measured from this simulation are $v_{\mathrm{travel}}=27.9\,\mathrm{km}\,\mathrm{s}^{-1},l_{\mathrm{apex}}=-49.1^{\circ},b_{\mathrm{apex}}=-80.4^{\circ}$; see the all-sky projection in Fig. 2 for the direction of the apex on-sky (pink cross).
@@
Motivated by the LMC mass posterior in Fig. 2 for the fiducial deforming MW–LMC simulation of Garavito-Camargo2019, we also take the other deforming simulations data from this suite that adopt the same MW model but have varying LMC masses i.e., $\mathrm{M_{LMC}}=\{8,10,\,25\}\times 10^{10}\,\mathrm{M}_{\odot}$. The LMC masses are defined to be their mass when each LMC first crossed the virial radius of the MW. Repeating the above inference methodology with these deforming MW–LMC simulations, we obtain the LMC mass posterior distribution when assuming knowledge of the LMC present day position, velocity, and reflex motion parameters. We compare the true simulation LMC mass to the median, plus $1\,\sigma$ confidence interval, of the SBI LMC mass posteriors in Fig. 3. For all except the most massive LMC model in Garavito-Camargo2019, we find that the true and SBI LMC masses are consistent, i.e., they overlap on the 1:1 grey line, within the $1\,\sigma$ confidence interval. To explain this, we re-weight the returned SBI median and percentiles to account for sampling the LMC mass from a Gaussian prior probability distribution and instead adopt a flat prior distribution (fainter points and error bars). This accounts for any biases that are introduced as a result of the choice of the Gaussian prior distribution. In Sec. 6.2 we discuss further the effect of the choice of prior. Now, all of the true and SBI LMC masses are consistent within their $1\,\sigma$ confidence intervals. Intuitively this agrees with expectation, as the most massive LMC, $M_{\rm LMC}=25\times 10^{10}\,\mathrm{M}_{\odot}$ is towards the lower probability tails of the adopted Gaussian prior distribution e.g., see grey contours in Fig. 2. Hence, any SBI estimate of the LMC mass would be expected to be biased to lower values based on the choice of our prior. Conversely, the opposite is also true for the least massive LMC; see Fig. 3. Furthermore, we demonstrate the expected increase in measured travel velocity (marker fill colour) as the LMC mass is increased (Petersen2021). Overall, our SBI framework is able to correctly recover the LMC mass well from more complex, deforming MW–LMC simulations. This is encouraging for future applications of this technique to constrain the real LMC mass using observational data (this has been performed in a simultaneous and distinct study, see Brooks2025).
Figure 3: Comparison of the true and inferred infall LMC masses: We compare the simulation truth and returned SBI posterior LMC mass for all of the MW–LMC deforming simulations in Garavito-Camargo2019. We show the $16^{\rm th}-84^{\rm th}$ percentiles of the LMC mass posteriors as error bars. For all deforming MW–LMC simulations except the most massive LMC, the true and SBI returned LMC masses are consistent, i.e., they overlap on the 1:1 dashed grey line, within the $1\,\sigma$ confidence interval. Additionally, we show the re-weighted SBI masses, which account for any biases in the choice of the LMC mass prior, as the fainter points and errors. We colour each point by the measured simulation travel velocity to demonstrate its expected increase with increasing LMC mass.